OpenStudy (anonymous):

Chapter 5:Indices and Logarithms @ganeshie8

OpenStudy (anonymous):

\[Given~\log _{3}2=0.631~\log _{3}5=1.465,find~the~values~of:\]\[\log _{3}1.2\]

ganeshie8 (ganeshie8):

Hint : \[\large 1.2 = \dfrac{12}{10} = \dfrac{6}{5} = \dfrac{2\times 3}{5}\]

OpenStudy (anonymous):

\[\log _{3}2+\log _{3}3-\log _{3}5\]

ganeshie8 (ganeshie8):

Excellent!

OpenStudy (anonymous):

The answer will be 0.166

OpenStudy (anonymous):

Thnx @ganeshie8

ganeshie8 (ganeshie8):

\[\large \color{Red}{\checkmark}\]